OpenStudy (anonymous):

Solve for p. (p ÷ 15) – 13 = 18 p = Solve for p. (p ÷ 15) – 13 = 18 p = @Mathematics

OpenStudy (anonymous):

add 13 to both sides then multiply both side by 15

OpenStudy (anonymous):

(p/15) = 18+13 p= (18+13) * 15 p = 465

OpenStudy (anonymous):

thanks